ganeshie8 (ganeshie8):

approximate \(\large \frac{2}{7}\) to 7 decimal places. Show all work for a gold medal

OpenStudy (kainui):

\[\Large \frac{2}{7} = 2 \frac{1}{1-(-6)}=2(1-6+6^2-6^3+ \cdots)\] Ok just kidding haha. =P

Parth (parthkohli):

long division...?

OpenStudy (anonymous):

Calculator ?

Parth (parthkohli):

just committed a plebian mistake

OpenStudy (anonymous):

oh okay

Parth (parthkohli):

did it on paper and it turns out \(0.\overline{285714}\)

OpenStudy (kainui):

I was going to guess something along the lines of this, but thought it would take too much work and didn't really amount to anything \[\Large \frac{2}{7} = \frac{1}{\frac{7}{2}}=\frac{1}{3+\frac{1}{2}}\]

Parth (parthkohli):

0.285714285714... that looks like it'd be 0.2857143 till seven digits.

OpenStudy (kainui):

Interesting and relevant http://primes.utm.edu/glossary/xpage/PeriodOfADecimal.html

OpenStudy (kainui):

So this says the period of 1/k is the same as the order of 10 mod k. Which means we have \[\Large 10^6 \equiv 1 \mod 7\] so 6 is the order and also the period. So it suffices that we will only have to compute this up to 6 decimal places instead of 10 I guess haha.

OpenStudy (abdullah1995):

this is some next level stuff lol

OpenStudy (xapproachesinfinity):

cool you have you (1+x)^r series expansion @Kainui

OpenStudy (xapproachesinfinity):

1-x i meant

OpenStudy (kainui):

What do you mean, like the geometric series or something else?

OpenStudy (kainui):

Honestly I don't even know why people are throwing medals at me, I didn't get this question right or anything lol

OpenStudy (xapproachesinfinity):

i meant how did you get that 2(1-6+6^2....) you did you use hehe

OpenStudy (xapproachesinfinity):

isn't that Tylor series stuff?

OpenStudy (kainui):

Yeah but if you calculate that it is clearly divergent. 1-6=5, 1-6+36=31, 1-6+36-6^3=-185, etc... I didn't and couldn't calculate any digits of the fraction with this.

OpenStudy (xapproachesinfinity):

oh i see i thought that would get you the digits well it is still cool though

OpenStudy (kainui):

Haha yeah whenever you can sum up a divergent series and get a value, this is called its Cesaro sum which is really quite an interesting thing since it's the limit of the average of the terms surprisingly.

OpenStudy (anonymous):

2/7 can be written as \[2\div7\], which you will get 2.0/7 = .2R .6, .60/7 = .08 R .04, .040/7 = .005R .005, .0050/7 = .0007R .0001, .00010/7 = .00001R .00003, .000030/7 = .000004R .000002, and wraps around. This is \[\approx .28571428\], so it would be rounded (in 7 decimal places) to .2857143